find how much water to add to a 360cm^3 of a 24% solution to give a 12% solution
Answer by mananth(16946)

-------------- percent ---------------- Amount
Solution 24 ---------------- 360 Cm^3
water 0 ---------------- x Cm^3
Mixture 12 ---------------- 360 + x Cm^3
Water conc is =0
24 * 360 + 0 x = 12 ( 360 + x )
8640 + 0 x = 4320 + 12 x
0 x -12 x = -8640 + 4320
-12 x = -4320
/ -12
x = 360 Cm^3 water

m.ananth@hotmail.ca